# Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences

### v1

@@ 3 Imbalances Analysis
In this section, we analyze the imbalance issues encountered during data sampling and packing in training on real-world datasets, and discuss potential optimization opportunities. To ease the analysis, we exemplify the imbalance issues by training the LLaMA 7B model with the CommonCrawl [78] and GitHub [16] datasets over 8 GPUs. The context length is 32K and the batch size is set as 100K for each iteration.
3.1 Data Sampling Imbalance
@@
Intra-iteration imbalance. The uneven sequence length distribution within an iteration can be attributed to the inherent property of data. In particular, sequences in the real world generally follow a long-tail distribution, where short sequences dominate and long sequences are rare. Existing studies have also observed such a distribution in many popular training datasets [19]. This imbalance poses both challenges and opportunities when it comes to optimizing overall performance.
@@
The challenges stem from the trade-offs between memory usage and throughput offered by different numbers of GPUs and parallel schemes. As shown in Figure 4, increasing the number of GPUs generally leads to a lower throughput per GPU but allows for a longer maximum supportable sequence length. Meanwhile, under the same number of GPUs, different parallel schemes (e.g., PP, TP and their combinations) also give divergent trade-offs. As aforementioned in §2.2, existing works utilize all GPUs in the cluster uniformly with a homogeneous parallel strategy tailored for the longest context length, compromising the training efficiency for memory.
@@
Nevertheless, opportunities arise when adapting the trade-offs to sequences of varying lengths: short sequences, which are more frequent, can be processed using faster but more memory-intensive parallel schemes, while the fewer long sequences can be assigned to slower yet more memory-efficient parallel schemes. Therefore, adopting heterogeneous parallel strategies— where non-unique parallel schemes are applied on different proportions of GPUs based on different sequence lengths— enables better performance (§5).
@@
Figure 4: Memory and throughput trade-offs. "OOM" indicates out-of-memory, and "N/A" indicates not available. The experiment is conducted using a 13B LLaMA model with 16 Nvidia A800 GPUs. Throughput is measured in tokens per second.
@@
Inter-iteration imbalance. While intra-iteration imbalance arises from the distribution of sequence lengths within a single mini-batch, another challenge comes from the distribution variations across mini-batches. Particularly, due to the long-tail distribution, rare long sequences may not appear in every iteration (even with a large batch size, the maximum sequence length still varies, as detailed in Appendix A.1). This leads to fluctuations in data distribution and maximum sequence lengths across iterations. However, current works overlook this issue, using a static parallel strategy for all iterations, which is unsuitable. For example, when an iteration consists mainly of short sequences, parallel schemes designed for long sequences become overkill. To address this, we propose to dynamically adjust the parallel strategy based on the characteristics of each mini-batch to better align the workload, thereby enhancing overall performance across iterations (§5).
Figure 5: The numbers of original sequences (classified by their lengths) in different micro-batches after packing (each micro-batch accounts for one packed sequence).
3.2 Data Packing Imbalance
@@
We then focus on the imbalance issues of data packing. As shown in Figure 5, we observe that due to varying sequence lengths, there are significant differences in the distributions of original sequences within the packed sequences, both among different micro-batches within the same pipeline and across different pipelines. In the following, we will analyze the impact of each aspect individually.
@@
Intra-pipeline imbalance. Since the memory consumption of each micro-batch is linear to the total length after packing, existing works commonly adopt fixed-length packing methods, like bin packing [86, 66], to balance memory usage. However, due to the quadratic time complexity of attention [80, 18, 66], the computation latency can hardly be balanced with fixed-length packing— for example with Figure 1 again, attention latency can be nearly 20×20$\times$20 × smaller when packing 32 sequences of 1K compared to a single 32K sequence.
@@
In pipeline parallel (PP), this leads to uneven execution times across micro-batches, yet common pipeline scheduling, such as 1F1B [33] or zero-bubble [69], requires uniform execution time of micro-batches to minimize pipeline bubbles. This highlights the importance of taking account of both the linear memory complexity and quadratic time complexity, rather than considering solely the memory as in fixed-length packing. To address this, we devise a nuanced packing strategy that considers the variation in original sequence lengths, striking a balance in terms of the execution time across different micro-batches within the memory constraint (§6.1).
@@
Inter-pipeline imbalance. Although different pipelines execute concurrently, the training latency is dictated by the slowest pipeline due to the need for gradients synchronization. Simply packing sequences to a fixed length has a similar issue of unbalanced computation load. In some cases, if the sequence number after packing cannot be evenly distributed across pipelines, a more severe imbalance will arise. In contrast to current systems that first pack sequences and then attempt to distribute them evenly, we suggest reversing this process. Specifically, our work first dispatches sequences of varying lengths to different pipelines, and then focuses on packing them within each pipeline. This innovative dispatching paradigm ensures a more balanced load across pipelines, before organizing sequences into micro-batches (§6.2).
4 Overview
@@
To tackle the data packing imbalance, given a mini-batch of sequences and a strategy, we devise a two-stage assignment (§6). The first stage determines how to dispatch these sequences of varying lengths to different pipelines to solve the inter-pipeline imbalance. Then, the second stage meticulously packs the sequences within each pipeline to form different micro-batches, addressing the intra-pipeline imbalance. This two-stage assignment process also estimates the latency of the given strategy to train the given mini-batch.
@@
5 System Designs
In this section, we start by mathematically defining our heterogeneous parallel strategies, then introduce the system techniques proposed to support these strategies and seamless transitions between them.
5.1 Dynamic Heterogeneous Strategies
@@
Next, we define our heterogeneous strategy space as the linear span of the parallel scheme space: $\mathbb{S}=\text{span}(\mathbb{P})$, meaning that each heterogeneous strategy can be represented as a linear combination of parallel schemes: $S=\sum_{i=1}^{K}d_{i}\times P_{i}$, which indicates that $S$ consists of $\sum_{i=1}^{K}d_{i}$ pipelines of $K$ different parallel schemes, with $d_{i}$ pipelines deployed with $P_{i}$.
@@
Strategy execution and transition. During each training iteration, given a mini-batch of sequences, we select the most suitable $S$ to execute. [^2] The selection in each iteration is based on the two-stage sequence assignment (§6), and we propose the candidate strategies by analyzing the sequence length distribution of the entire dataset (§7). The pipelines will be assigned with non-overlapping sequences and execute the forward and backward propagation concurrently, which is termed the propagation phase. Then, akin to traditional data parallel (DP), all pipelines synchronize the model gradients and perform updates, which is termed the optimization phase. Since different iterations inevitably select non-unique strategies, strategy transitions between iterations are necessary.
System challenges. However, enabling the system to support various pipelines, each configured with a different $P$, is complex due to the need to manage diverse communication patterns across devices. Furthermore, dynamically transitioning between strategies at the iteration level presents additional difficulties, as the model states mappings on devices can vary significantly across different $S$. To address these issues, we develop an optimization-propagation disaggregation approach (§5.2) and a subgraph-based abstraction (§5.3). These system designs successfully meet our requirements of training with dynamic, heterogeneous strategies.
5.2 Optimization-propagation Disaggregation
@@
As illustrated in Figure 2(b), unlike traditional methods that employ a uniform distributed training strategy, we separate the optimization phase from the propagation phase, decoupling the GPU’s role as a distributed storage from its computational role during propagation. Specifically, this approach maintains a constant partitioning for FP32 parameters and optimizer states, while allowing any parallel strategy to be used during propagation, enabling the system to support dynamic heterogeneous strategies and seamless transitions between them. [^3] We use a fully sharded strategy in optimization phase, yet any other zero-redundancy sharding strategies are also feasible. When propagation requires access to FP32 parameters and optimizer states, we perform pull and push interactions, which correspond to the retrieval of parameters and the synchronization and scattering of gradients, respectively. These interactions occur only at the beginning and end of a training iteration, referred to the steps 1-2 and steps 4-5 of Figure 2(b).
To effectively manage pull and push interactions, we develop a specialized LNK (link) operator to handle complex interaction patterns. As depicted on the right side of Figure 7, this operator utilizes three core primitives: send, receive, and local movement, centered around a fundamental unit called the mutual slice, which represents the common partitioning granularity when model shardings interact between the propagation and optimization phases. Given that the optimization states are partitioned on each GPU without redundancy, the pull and push schemes can be uniquely determined using these three primitives on the non-duplicate mutual slices.
In particular, assume there are $N$ GPUs, the optimization phase employs a fully sharded strategy, and the propagation phase uses a strategy $S=\sum_{i=1}^{K}d_{i}\times P_{i}$ with $D=\sum_{i=1}^{K}d_{i}$ pipelines. Let $\overline{TP}_{\text{max}}=\max_{i\in[1,K]}\texttt{TP}(P_{i})$. For the pull interaction, the mutual slices are created by evenly partitioning the global parameter into $\mathtt{LCM}\left(N,\overline{TP}_{\text{max}}\right)$ segments (where $\mathtt{LCM}$ denotes the least common multiple). Each GPU identifies its necessary mutual slices. If a slice is local, it uses the local movement operation; otherwise, it employs send and receive to exchange slices. Similarly, for the push interaction, as shown on the left side of Figure 7, a reduce-scatter is first performed at the corresponding location to synchronize the gradient. Then the partitioning granularity of the mutual slice is $\mathtt{LCM}\left(N,D\times\overline{TP}_{\text{max}}\right)$, with the additional $D$ accounting for the gradient’s reduce-scatter. The LNK can be established still using the send, receive and local movement primitives.
@@
Figure 7: Diagram of RS (reduce-scatter) and LNK (link) operations during push. Pink indicates the gradient, blue the optimizer states, and green the mutual slice. The optimizer states are partitioned with a granularity of 4, and the gradient after reduce-scatter has a granularity of 6. Thus, the mutual slice has a granularity of 12 (i.e., the least common multiple of 4 and 6).
@@
Compared to training with homogeneous strategies, introducing the LNK operator incurs additional overhead to resolve the discrepancy between strategies. However, the send and receive communication can be overlapped with computation during propagation, and the local movement incurs only minimal latency, so the overhead is small (as evaluated in §8.4). Last but not least, our design offers great generalization capabilities— when the optimization and propagation use the same strategy, the LNK operator in the pull will be translated into an all-gather operation, and the LNK operator in the push will be pruned (leaving a reduce-scatter operation).
5.3 Subgraphs

### v2

@@ 3. Imbalances Analysis
In this section, we analyze the imbalance issues encountered during data sampling and packing in training on real-world datasets, and discuss potential optimization opportunities. To ease the analysis, we exemplify the imbalance issues by training the LLaMA 13B model with the CommonCrawl (TIIUAE, [n. d.]) and GitHub (CodeParrot, [n. d.]) datasets over 16 GPUs, with 32K context length and 100K tokens per iteration.
Figure 4. The dot plot on the left shows the fluctuation of maximum sequence length, emphasizing inter-iteration imbalance, while the bar chart on the right presents the distribution of sequence lengths within an iteration (20th iteration), highlighting intra-iteration imbalance.
3.1. Data Sampling Imbalance
@@
The uneven sequence length distribution within an iteration can be attributed to the inherent property of data. In particular, sequences in the real world generally follow a long-tail distribution, where short sequences dominate and long sequences are rare. Existing studies have also observed such a distribution in many popular training datasets (de Vries, 2023). This imbalance poses both challenges and opportunities when it comes to optimizing overall performance.
@@
The challenges stem from the trade-offs between memory usage and throughput offered by different numbers of GPUs and parallel schemes. As shown in Figure 5, increasing the number of GPUs generally leads to a lower throughput per GPU but allows for a longer maximum sequence length. Meanwhile, under the same number of GPUs, different parallel schemes (e.g., PP, TP and their combinations) also have divergent trade-offs (e.g., replacing PP with TP generally supports longer sequences but at the cost of reduced per-GPU throughput). As aforementioned in §2.2, existing works utilize all GPUs in the cluster uniformly with a homogeneous parallel strategy tailored for the longest context length, compromising the overall training efficiency for memory.
@@
Nevertheless, combined with the length imbalance within a mini-batch, such trade-offs become particularly invaluable: short sequences can be processed using faster but more memory-intensive parallel schemes, while the long sequences can be assigned to slower yet more memory-efficient parallel schemes. Therefore, adopting heterogeneous parallel strategies— where non-unique parallel schemes are applied on different proportions of GPUs based on different sequence lengths— enables better performance (§5).
@@
Figure 5. Memory and throughput trade-offs. “OOM” indicates out-of-memory, and “N/A” indicates not available. The experiment is conducted using a 13B LLaMA model with 16 Nvidia A800 GPUs. Throughput is measured in tokens per second and maximum sequence lengths are given by $\texttt{MaxLen}(\cdot)$ (see §6.1 and Appendix C.1 (hyd, 2025)).
Inter-iteration imbalance
@@
While intra-iteration imbalance arises from the distribution of sequence lengths within a single mini-batch, another challenge comes from the distribution variations across mini-batches. Particularly, due to the long-tail distribution, rare long sequences may not appear in every iteration (even with a large batch size on a 1024-GPU cluster, the maximum sequence length still varies, as detailed in Appendix A.1 (hyd, 2025)). This leads to fluctuations in data distribution and maximum sequence lengths across iterations. However, current works overlook this issue, using a static parallel strategy for all iterations, which is unsuitable. For example, when an iteration consists mainly of short sequences, parallel schemes designed for long sequences become overkill. To address this, we propose to dynamically adjust the parallel strategy based on the characteristics of each mini-batch to better align the workload, thereby enhancing overall performance across iterations (§5).
Discussions
[…]
Additionally, a prior work, HotSPa (Ge et al., 2024), addresses the data sampling imbalance in a different way—it performs length-based training within each iteration instead of over the entire dataset, in order to preserve the model convergence. By dynamically switching between expert-tuned, homogeneous parallel strategies for different sequence lengths, HotSPa accumulates gradients and performs a single global update per iteration. However, it lacks support for heterogeneous strategies and finer-grained data assignment, and frequent strategy switching introduces additional overhead. We will empirically compare HotSPa and our work in §8.
Figure 6. Training and evaluation loss of different sampling methods. We subsample 250M tokens for training and 5M for validation from the GitHub dataset, and train a 13B model on 16 GPUs with 32K context length and 100K tokens per iteration.
3.2. Data Packing Imbalance
@@
We then focus on the imbalance issues of data packing. As illustrated in Figure 7, we observe significant distribution differences in the original sequences within packed sequences (i.e., micro-batches) due to varying sequence lengths. These variations occur both within micro-batches of the same pipeline and across different pipelines, resulting in execution time variance. Below, we analyze each aspect’s impact separately.
Intra-pipeline imbalance
@@
Since the memory consumption of each micro-batch is linear to the total length after packing, existing works commonly adopt max-length packing methods, like bin packing (Yao, 1980; NVIDIA, 2023d), to balance memory usage. However, due to the quadratic time complexity of attention (Vaswani et al., 2017; Dao et al., 2022; NVIDIA, 2023d), the computation latency can hardly be balanced in this way— for example with Figure 2 again, attention latency can be nearly $20\times$ smaller when packing 32 sequences of 1K compared to a single 32K sequence.
@@
In pipeline parallel (PP), this leads to uneven execution times across micro-batches, yet common pipeline scheduling, such as 1F1B (Harlap et al., 2018) or zero-bubble (Qi et al., 2023), requires uniform execution time of micro-batches to minimize pipeline bubbles. This highlights the importance of taking account of both the linear memory complexity and quadratic time complexity, rather than considering solely the memory as in max-length packing. To address this, we devise a nuanced packing strategy that considers the variation in original sequence lengths, striking a balance in terms of the execution time across different micro-batches within the memory constraint (§6.1).
Inter-pipeline imbalance
@@
Although different pipelines execute concurrently, the training latency is dictated by the slowest pipeline due to the need for gradient synchronization. Simply packing sequences to a fixed maximum length has a similar issue of unbalanced computation load. In some cases, if the sequence number after packing cannot be evenly distributed across pipelines, a more severe imbalance will arise. In contrast to current systems that first pack sequences and then attempt to distribute them evenly, we suggest reversing this process. Specifically, our work first dispatches sequences of varying lengths to different pipelines, and then focuses on packing them within each pipeline. This innovative dispatching paradigm ensures a more balanced load across pipelines, before organizing sequences into micro-batches (§6.2).
Figure 7. The numbers of original sequences (classified by their lengths) in different micro-batches after packing (each micro-batch accounts for one packed sequence) and their corresponding time cost. The experiment is conducted using a 13B LLaMA model with 16 Nvidia A800 GPUs.
4. Overview
@@
To tackle the data packing imbalance, given a mini-batch of sequences and a strategy, we devise a two-stage assignment (§6). The first stage determines how to dispatch sequences of varying lengths to different pipelines to solve the inter-pipeline imbalance. Then, the second stage meticulously packs the sequences within each pipeline to form different micro-batches, addressing the intra-pipeline imbalance. This two-stage assignment process also estimates the latency of the given strategy to train the given mini-batch.
@@
5. System Designs
In this section, we start by mathematically defining our heterogeneous parallel strategies, then introduce the system techniques proposed to support these strategies and the dynamic transitions between them.
5.1. Dynamic Heterogeneous Strategy Design
Definitions
@@ 5.1. Dynamic Heterogeneous Strategy Design
$\mathbb{P}=\big\{P|$ $\texttt{TP}(P)\in\mathbb{D}_{T},\texttt{PP}(P)\in\mathbb{D}_{P},\texttt{CP}(P)\in\mathbb{D}_{C},$
$\texttt{TP}(P)\times\texttt{PP}(P)\times\texttt{CP}(P)\leq N\big\},$
where $\mathbb{D}_{T}=\mathbb{D}_{C}=\{1,2,\cdots,N\}$, $\mathbb{D}_{P}=\left\{v\in\mathbb{N}^{+}|v\text{ divides }\overline{\textup{layers}}\right\}$, $N$ is the number of GPUs, and $\overline{\textup{layers}}$ is the number of layers.
@@
Next, we define our heterogeneous strategy space as the linear span of the parallel scheme space: $\mathbb{S}=\text{span}(\mathbb{P})$, meaning that each heterogeneous strategy can be represented as a linear combination of parallel schemes: $S=\sum_{i=1}^{K}d_{i}\times P_{i}$, which indicates that $S$ consists of $\sum_{i=1}^{K}d_{i}$ pipelines of $K$ different parallel schemes, with $d_{i}$ pipelines deployed with $P_{i}$. Figure 8 shows some example parallel schemes ($P$) and heterogeneous strategies ($S$) when the number of GPUs ($N$) is 8.
Strategy execution and transition
@@
During each training iteration, given a mini-batch of sequences, we select the most suitable $S$ to execute. [^3] The selection in each iteration is based on the two-stage sequence assignment (§6), and we propose the candidate strategies by analyzing the sequence length distribution of the entire dataset (§7). The pipelines will be assigned different sequences and execute forward and backward passes concurrently to compute gradients from the corresponding parameters. This process is termed the propagation phase (Figure 3(a)3(b), step 3). Then, similar to traditional data parallelism (DP), all pipelines synchronize the model gradients and perform updates at the end of each iteration to obtain the new parameters and optimizer states. This process is called the optimization phase (Figure 3(a)3(b), step 6). Since different iterations inevitably select non-unique strategies, strategy transitions between iterations are necessary (e.g., in Figure 8, iteration 8 uses strategy $S5$, while iteration 9 transitions to strategy $S2$).
Challenges
To support the execution and transition of arbitrary heterogeneous strategies, the conventional approach to handling model states must be revised. As shown in Figure 3(a), for current systems that employ a static, homogeneous parallel strategy, it is common practice to shard the model parameters and optimizer states according to the employed homogeneous strategy, and leverage all-gather and reduce-scatter operations to retrieve the parameters and synchronize the gradients, respectively. However, this approach falls short in supporting dynamic heterogeneous strategies, as model partitioning varies significantly across pipelines. For example, in Figure 3(b), pipelines with different TP degrees produce parameters and gradients with divergent partitioning granularities (i.e., tensors are split into non-uniform slices across GPUs). Consequently, it is infeasible to discover a single communication pattern to accomplish the parameter retrieval or gradient synchronization. This limitation drives our development of the LNK (link) operator, which supports parameter retrieval and gradient synchronization in an optimization-propagation disaggregated manner (§5.2), and we further introduce subgraphs to encapsulate diverse communication patterns (§5.3). Furthermore, since gradient synchronization requires all pipelines to participate and is constrained by the slowest one, pipeline heterogeneity also poses significant load-balancing challenges, which we will address in §6.
5.2. Optimization-Propagation Disaggregation
@@
As illustrated in Figure 3(b), we separate the optimization phase from the propagation phase, decoupling the GPU’s role as a distributed storage from its computational role during propagation. Specifically, we keep the partitioning of FP32 parameters and optimizer states fixed for the optimization phase [^4] We use a fully sharded strategy in the optimization phase, yet any other zero-redundancy sharding strategies are also feasible., while allowing an arbitrary parallel strategy to be used in the propagation phase. In such a disaggregated design, to ensure correct parameter partitioning during propagation as well as correct gradient partitioning for optimizer update, we establish two types of interaction: (a) pull, which retrieves the parameters from the optimization phase to an arbitrary heterogeneous parallel strategy (Figure 3(b), steps 1–2); (b) push, which synchronizes and scatters gradients back to the optimization phase (Figure 3(b), steps 4-5). Compared to the all-gather and reduce-scatter in the conventional design that requires identical, homogeneous parallel strategy in the two phases (Figure 3(a), steps 1-2 and 4-5), our pull and push abstractions enable the use of arbitrary heterogeneous parallel strategies in the propagation phase as well as the seamless transition between different strategies.
@@
Figure 9. Diagram of RS (reduce-scatter) and LNK (link) operations during push. Pink and blue denote slices in the propagation and optimization phases, respectively, while green represents the mutual slice (the finest-grained slice shared by both). The partitioning granularity is 4 in the propagation phase and 6 in the optimization phase (after reduce-scatter). Thus, the mutual slice has a granularity of 12 (i.e., the least common multiple of 4 and 6).
To effectively manage pull and push interactions, we develop a specialized LNK (link) operator to handle complex communication patterns. As depicted on the right side of Figure 9, this operator utilizes three core primitives: send, receive, and local movement, all operating on a fundamental unit called the mutual slice, which represents the finest-grained data partition. These primitives enable dynamic resharding of parameters and gradients between propagation and optimization phases. For instance, in Figure 9, GPU1 generates a gradient slice during propagation, which is further partitioned into two finest-grained mutual slices (7 and 8). Since GPU0 requires mutual slice 7 during the optimization phase for model updates, it triggers a send to transfer mutual slice 7 from GPU1 to GPU0. Because optimization states are uniquely partitioned across GPUs (without redundancy), the use of these primitives is deterministic.
Mathematically, assume there are $N$ GPUs, the optimization phase employs a fully sharded strategy, and the propagation phase uses a strategy $S=\sum_{i=1}^{K}d_{i}\times P_{i}$. Let $\overline{TP}_{\text{max}}=\max_{i\in[1,K]}\texttt{TP}(P_{i})$. For the pull interaction, the mutual slices are created by evenly partitioning the global parameter into $\mathtt{LCM}\left(N,\overline{TP}_{\text{max}}\right)$ segments (where $\mathtt{LCM}$ denotes the least common multiple). Each GPU identifies its necessary mutual slices. If a slice is local, it uses the local movement operation; otherwise, it employs send and receive to exchange slices. Similarly, for the push interaction, as shown on the left side of Figure 9, a reduce-scatter is first performed at the corresponding location to synchronize the gradient. Then the partitioning granularity of the mutual slice is $\mathtt{LCM}\left(N,D\times\overline{TP}_{\text{max}}\right)$, where the additional factor $D=\sum_{i=1}^{K}d_{i}\times\texttt{CP}(P_{i})$ accounts for gradients reduce-scatter across all model replicas. The LNK can be established still using the send, receive and local movement primitives. Notably, the LNK mechanism operates independently of $\texttt{PP}(P_{i})$, as PP doesn’t partition individual parameters into slices. From the perspective of heterogeneous strategy execution, the LNK operator resides in different stages across training pipelines. Consequently, communication occurs only after the slowest pipeline reaches the LNK operator (the balancing across pipelines will be addressed in §6).
@@
It is noteworthy that our optimization-propagation disaggregation is fundamentally a generalization of current model states sharding techniques (i.e., ZeRO (Rajbhandari et al., 2020) and FSDP (Zhao et al., 2023))— when using homogeneous strategies, the LNK operator in the pull will be translated into an all-gather operation, and the LNK operator in the push will be pruned (leaving a reduce-scatter operation), which are the same as current techniques. For heterogeneous strategies, introducing the LNK operator incurs additional overhead to resolve the discrepancy between strategies. However, the send and receive communication can be overlapped with computation during propagation, and the local movement incurs only minimal latency, so the overhead is small (as evaluated in §8.4).
Last but not least, theoretically speaking, the communication during pull and push operations is bounded by the model size and does not scale with the number of GPUs. And the worst-case communication volume is exactly the same as that of current model states sharding techniques. Furthermore, computation-communication overlap can be achieved when each micro-batch in the pipeline with $P_{i}$ contains more than $\texttt{TP}(P_{i})\times\texttt{CP}(P_{i})\times\frac{F}{B}$ tokens, where $F$ is the GPU’s half-precision compute capability (FLOPS), and $B$ is the slowest interconnect bandwidth (bytes/sec). This condition is independent of the model size or the number of GPUs. Due to the space constraint, detailed proof and analysis are provided in Appendix B (hyd, 2025).
5.3. Subgraphs